A Complete Bibliography of Publications in Aquaculture, Fish and Fisheries

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

28 December 2023 Version 1.13

Title word cross-reference

× [FFH^{+}22, LDJL^{+}24].

//doi.org/10.1002/aff2.81 [Ano23b].

19 [ADY^{+}22].

2023 [ZOB^{+}23].

A. [LDJL^{+}24]. abundance [RSAM23]. acid [HPPL23, HMK^{+}22, OOMM23, STF^{+}22]. acidification [UOCT23]. Acipenser [CMR23, NG23]. activity [RZA^{+}22]. acute [DJF^{+}22]. adjacent [UvdRJ23]. adult [LDJL^{+}24]. adults [DJF^{+}22]. Aeromonas [PFM^{+}22]. aff2.81 [Ano23b]. affect [CTCH23, LDJL^{+}24, UOCT23]. Africa [ADY^{+}22, BDGS21, MVN^{+}24]. African [AKK^{+}23, KNM^{+}23, OYO22, SPZT23]. after [RCC^{+}23, STF^{+}22]. age [BWM22]. age-frequency [BWM22]. algal [OOMM23, SPZT23]. alterations [DJF^{+}22, UPSP22]. alternative [AMS^{+}22]. Amaranthus [MBT^{+}22]. amazonicum [DJF^{+}22]. ammonia [DJF^{+}22]. among [KG23]. Anabas [RDH^{+}22]. analysis [BDGS21, DTY^{+}21, IM22, NG23, OYO22, RBW^{+}22, SLU^{+}23, TAL^{+}22]. Anarhichas [LDJL^{+}24]. anchovy [NIYK22]. antioxidant [RZA^{+}22]. apoptosis [UPSP22]. applied [HP22]. Appraisal [DAJ^{+}23]. approach [ZOB^{+}23]. approaches [HBP22]. Aquaculture [Ano23b, PC21, ZOB^{+}23, APT^{+}22, ADY^{+}22, BDGS21, LL23, MBD^{+}23, OMO^{+}23, RBW^{+}22, SLU^{+}23, USB^{+}22, BC21]. aquaculture-related [ADY^{+}22]. aquaponic [SMM^{+}22]. aquarium [ASS^{+}23]. aquatic [MVN^{+}24, RZA^{+}22]. aquatica [RZA^{+}22]. Arapaima [RCC^{+}23]. Arctic [AOR^{+}22]. arcticus [AOR^{+}22]. area [RCC^{+}23, SMD^{+}22]. Artemia [OOMM23]. artificial [KTT22]. artisanal [MMG23]. aspects [SBJJ22]. associations [GFL22]. atkinsoni [TDT23]. Atlantic [ASS^{+}23, BWM22, PGT^{+}23, RPM^{+}23, RBW^{+}22]. attached [MLHLGV^{+}23]. attitudes [KG23]. auratus [UvdRJ23]. Australia [CMN^{+}23]. autochthonous [PFM^{+}22]. Automated [TAL^{+}22, WNS^{+}22]. availability [AOR^{+}22, Pla22].

B [Ano23b]. backcrossbreds [RDH^{+}22]. Backpack [Ano23b, TE22]. bacteria [CMW^{+}22, KOK^{+}22, PFM^{+}22]. bacterial [HMS^{+}23]. Bangladesh [ARJ^{+}22, SAW^{+}22]. Bangladeshi [RDH^{+}22]. based [DAJ^{+}23, MKPL23]. Basin [MMG23, MMG^{+}22]. bata [ASSR21, ASSR21]. batrachus [PFM^{+}22]. Bayesian [RPM^{+}23]. beaver [AOR^{+}22]. behaviour [CMR23, KSH23, NF22]. behavioural [ASS^{+}23]. Benin [ACA22, ADY^{+}22, KNM^{+}23]. benthic [BAL^{+}22]. betta [MEHS22, MEHS22]. between [NIYK22]. Bight [BWM22]. bimaculatus [IHA^{+}22]. Bio [ADM23]. Bio-ecological [ADM23]. bioactive [RZA^{+}22]. biochemistry [DFG22]. Biofloc [MBD^{+}23, OOMM23, ZOGO22, IIB^{+}23]. biology [HBP22, MMG23, PC21]. biomarkers [LDJL^{+}24]. biomass [IM22]. Biotic [CMR23]. bivalve [TLS^{+}23]. Black [LSU^{+}22]. blue [LMD^{+}23, BC21]. boiling [STF^{+}22]. Bolivia [RCC^{+}23]. border [RCC^{+}23]. bottom [OYO22]. Brachionus [KSH23]. Branchitis [CMW^{+}22]. Brassica [MBT^{+}22]. breeding [RDH^{+}22]. Brimsu [ADM23]. brown [STF^{+}22]. bubble [Ano23b, TE22]. Buccinum [BWM22]. built [MMG^{+}22]. Burchell [SPZT23]. butter [IHA^{+}22].

Côte [AKK^{+}23]. cage [MKPL23, NF22, OMOO22]. Cameroon [SEN23]. canadensis [FFH^{+}22]. cannibalism [KTT22]. captivity [ERF^{+}23, Pla22]. carcass [BAL^{+}22]. case [ASS^{+}23, MLHLGV^{+}23, OMO^{+}23]. catch [Los22]. catfish [BSH^{+}22, HHR^{+}23, IHA^{+}22, SPZT23]. caught [SBJJ22]. cause [ASS^{+}23]. causes [BAL^{+}22]. cell [UPSP22]. Census [GRJM^{+}23]. cephalopod [EAAGCMR23]. chain [PGT^{+}23]. chain-1 [PGT^{+}23]. chains [SLU^{+}23]. challenge [PFM^{+}22]. challenges [MMG^{+}22]. Changes [STF^{+}22, ASS^{+}23, ERF^{+}23]. Characterisation [MNN^{+}22]. characteristics [TDT23]. characterization [ARJ^{+}22]. chilling [MKZ^{+}21]. Chinook [HP22, Los22, TAL^{+}22]. chlorination [SJ23]. chromatophoroma [MEHS22]. chronic [LDJL^{+}24]. Chrysophrys [UvdRJ23]. cichlids [ADM23]. clam [HPPL23]. Clarias [PFM^{+}22, SPZT23]. climbing [RDH^{+}22]. clutch [HBP22]. coastal [KNM^{+}23, RSAM23]. cod [ASS^{+}23]. coleoid [EAAGCMR23]. Colorado [CMW^{+}22]. combined [DJF^{+}22]. Combining [HBP22, Kit22, DCM^{+}23]. commercial [KOK^{+}22]. commercialised [UOCT23]. Commonwealth [TDT23]. communities [GFL22]. community [KG23, SPZT23]. Comparative [OYO22, SBJJ22, SPZT23]. Comparison [BDM21]. competitors [GRJM^{+}23]. Composition [NG23, MMM21, OOMM23, STF^{+}22]. compounds [RZA^{+}22]. computer [DCM^{+}23]. concentrate [NHP^{+}22]. concept [HBP22]. condition [DFG22, ERF^{+}23, MKZ^{+}21, NHP^{+}22, OOMM23, Pla22]. conditions [OOMM23]. constituents [HMS^{+}23]. contents [NG23]. contribute [Ano23b, TE22]. copepod [KTT22]. core [HMS^{+}23]. corn [NHP^{+}22]. Correspondence [Ano23b]. Corrigendum [Ano23a]. counts [NHP^{+}22]. COVID [ADY^{+}22]. COVID-19 [ADY^{+}22]. crab [SAW^{+}22]. Crangon [STF^{+}22]. Crassostrea [OYO22]. critical [SLU^{+}23]. cross [KG23]. cross-sectional [KG23]. crossbreds [RDH^{+}22]. crude [KMMM23]. cultivated [OYO22]. culture [DAJ^{+}23, GRJM^{+}23, MNI^{+}22, OMOO22, OYO22]. culture-based [DAJ^{+}23]. cultured [OOMM23]. Culturing [KOK^{+}22]. current [Kit22, PC21]. curve [EAAGCMR23]. cycle [HP22, UOCT23, UPSP22]. cyprinid [DTY^{+}21].

Dam [MMM21]. data [BWM22]. decisions [SEN23]. degraded [BAD^{+}22]. dehydrogenase [PGT^{+}23]. Delta [SMD^{+}22, AKH^{+}23, ERF^{+}23, CTCH23]. density [KTT22, SMM^{+}22, ZJS^{+}21]. Densu [OYO22]. deposition [BAL^{+}22]. detection [PGT^{+}23]. Development [RSAM23, TLS^{+}23, ZOB^{+}23, OMO^{+}23, SLU^{+}23]. diagnostic [PGT^{+}23]. Diagnostics [CMW^{+}22]. diel [CTCH23]. Diet [FFH^{+}22, UvdRJ23, KOK^{+}22]. dietary [PFM^{+}22]. diets [APT^{+}22, HPPL23]. different [AKJ^{+}23]. digestibility [KMMM23]. diploid [ASSR21]. Diplostomum [USB^{+}22]. disruptive [ADY^{+}22]. dissected [MLHLGV^{+}23]. distribution [LMD^{+}23]. distributional [ASS^{+}23]. distributions [RSAM23]. diversity [ACA22, MMM21, MMG^{+}22]. d’Ivoire [AKK^{+}23]. DNA [IM22]. Does [LDJL^{+}24, Ano23b, TE22]. doi.org [Ano23b]. domesticated [MKZ^{+}21]. drift [CMR23]. droebachiensis [SBJJ22]. Duration [MKZ^{+}21]. during [GRJM^{+}23]. dye [AKH^{+}23]. dynamics [SAW^{+}22]. earthen [AKJ^{+}23]. East [MVN^{+}24]. ecological [ADM23]. ecology [AKK^{+}23, LMD^{+}23, PC21]. Economic [ARJ^{+}22, OMOO22, ACA22, ADY^{+}22, AKK^{+}23, LSU^{+}22, MNI^{+}22]. economics [IIB^{+}23]. ecosystem [MVN^{+}24, ZOB^{+}23]. ecosystems [AKK^{+}23]. Ecuador [BDM21]. edulis [ZJS^{+}21]. Edward [MVN^{+}24]. edwardsii [DFG22]. Effect [IHA^{+}22, IIB^{+}23, NHP^{+}22, RPM^{+}23, ZJS^{+}21]. effectors [WNS^{+}22]. Effects [BSH^{+}22, HMK^{+}22, NYN^{+}22, SMM^{+}22, BAL^{+}22]. Efficacy [HHR^{+}23, AMS^{+}22]. efficiency [LSU^{+}22, MBT^{+}22, ZOGO22]. egg [NYN^{+}22]. electrofishing [Ano23b, TE22]. Elmina [KG23]. Emerging [MNI^{+}22]. emperor [TDT23]. Emphasis [USB^{+}22]. emptying [RPM^{+}23]. encapsulated [MLHLGV^{+}23]. endangered [ERF^{+}23]. Engraulis [NIYK22]. enhanced [SBJJ22]. enzyme [KMMM23]. Esox [GFL22]. establishment [MMG^{+}22]. estimated [RPM^{+}23]. estimates [CMN^{+}23]. estuarine [RSAM23]. Estuary [OYO22]. Ethiopia [MMG^{+}22]. ethyl [UPSP22]. European [ZJS^{+}21]. euryhaline [KSH23]. Evaluation [BWM22, PFM^{+}22, RDH^{+}22, SJ23, MKPL23, TLS^{+}23]. evaluations [AKH^{+}23]. evidence [ZOGO22]. expansion [MBD^{+}23]. Experimental [ZOGO22]. explains [Kit22]. exposed [CMW^{+}22]. exposure [DJF^{+}22, LDJL^{+}24]. extensive [SMD^{+}22]. external [Ano23b, TE22].

F [Ano23b]. faced [AKK^{+}23]. facilitates [RBW^{+}22]. factor [ERF^{+}23]. factors [SLU^{+}23, SEN23]. farm [AKJ^{+}23, ACA22]. farmed [BDM21, HMK^{+}22]. farming [ACA22, ARJ^{+}22, BAD^{+}22, MKPL23, NF22, SPZT23, SMD^{+}22]. farms [UvdRJ23]. fatty [HPPL23, HMK^{+}22, OOMM23, STF^{+}22]. feasibility [OMOO22, SMD^{+}22]. fecundity [ERF^{+}23]. fed [HPPL23, KMMM23]. feed [BSH^{+}22, HHR^{+}23, IIB^{+}23, LSU^{+}22, SLU^{+}23, WNS^{+}22]. feeders [WNS^{+}22]. female [MKZ^{+}21]. fermented [NHP^{+}22]. financial [SMD^{+}22]. Fish [ACA22, Ano23b, BC21, BAL^{+}22, MMM21, SLU^{+}23, APT^{+}22, AKK^{+}23, BSH^{+}22, IM22, KOK^{+}22, MKPL23, MNN^{+}22, MEHS22, NF22, NYN^{+}22, OMOO22, RSAM23, ACA22]. fished [Kit22]. fisher [AKK^{+}23]. Fisheries [Ano23b, RCC^{+}23, DAJ^{+}23, IM22, MMG^{+}22, MVN^{+}24, OMO^{+}23, SEN23, BC21]. fishers [KG23]. fishery [Los22, MMG^{+}22, PC21]. fishes [AOR^{+}22, HBP22]. fishing [KG23, MMG23]. fishmeal [HHR^{+}23]. fishmeal-replacer [HHR^{+}23]. fitting [EAAGCMR23]. flat [ZJS^{+}21]. flavescens [FSS^{+}23]. floating [GRJM^{+}23]. Florida [RSAM23]. flow [USB^{+}22]. flow-through [USB^{+}22]. fly [LSU^{+}22]. foods [KOK^{+}22]. formalin [NYN^{+}22]. formulated [HHR^{+}23]. formulation [TLS^{+}23]. fossilis [BSH^{+}22, HHR^{+}23, PFM^{+}22]. framework [ZOB^{+}23]. franciscana [OOMM23]. frequency [BWM22]. freshwater [AKK^{+}23, DAJ^{+}23, IIB^{+}23, KNM^{+}23]. fry [BSH^{+}22, LSU^{+}22, PC21, ZOGO22]. fulvescens [NG23]. function [EAAGCMR23]. gariepinus [SPZT23]. gas [Ano23b, TE22]. gastrointestinal [NG23]. gastropod [BWM22]. gene [PGT^{+}23]. general [EAAGCMR23]. generated [KOK^{+}22]. generations [ERF^{+}23]. Genetic [DTY^{+}21, Kit22, MNN^{+}22]. genotoxicity [UPSP22]. Geographical [KNM^{+}23]. Germany [USB^{+}22]. Ghana [ZOB^{+}23, ADM23, KG23, OYO22]. giant [DAJ^{+}23, IIB^{+}23]. gigas [RCC^{+}23]. Gill [DJF^{+}22, IHA^{+}22]. GIS [MKPL23]. GIS-based [MKPL23]. Glaucosoma [CMN^{+}23]. goby [PC21]. goby-fry [PC21]. Gompertz [HP22]. gonad [MKZ^{+}21]. gonads [SBJJ22]. granulomas [MLHLGV^{+}23]. grayling [AOR^{+}22]. green [OOMM23, SBJJ22, UvdRJ23]. green-lipped [UvdRJ23]. grounds [NIYK22]. group [CTCH23]. Growth [ASSR21, HPPL23, AKJ^{+}23, BSH^{+}22, CMN^{+}23, EAAGCMR23, FFH^{+}22, LDJL^{+}24, IHA^{+}22, IIB^{+}23, KMMM23, LSU^{+}22, NHP^{+}22, OYO22, PFM^{+}22, RDH^{+}22, TAL^{+}22, ZOGO22, ZJS^{+}21]. Gulf [MKPL23]. gut [GFL22].

habitat [AOR^{+}22, BAD^{+}22, FSS^{+}23]. habitats [UvdRJ23]. haematological [DJF^{+}22, IHA^{+}22]. haemocyte [NHP^{+}22]. haemolymph [DFG22]. Hamilton [ASSR21]. hapa} [GRJM^{+}23]. hard [HPPL23]. harvest [DFG22, Los22]. hatcheries [TLS^{+}23]. hatchery [CMW^{+}22, Los22]. Hawaiian [APT^{+}22]. health [BSH^{+}22]. helleri [HMK^{+}22]. hepatopancreas [NHP^{+}22]. Herbicide [UPSP22]. Herbst [SAW^{+}22]. Herklots [KNM^{+}23]. Hermetia [LSU^{+}22]. Heteropneustes [BSH^{+}22, HHR^{+}23, PFM^{+}22]. Hippocampus [Pla22]. histological [HBP22, NHP^{+}22]. histomorphology [IHA^{+}22]. history [TDT23]. HIV [KG23]. holdfast [Pla22]. Holothuria [GRJM^{+}23]. home [BC21]. https [Ano23b]. hybrid [LDJL^{+}24]. hybridus [MBT^{+}22]. Hydrocynus [MMG23]. hydrosystems [KNM^{+}23]. Hypomesus [ERF^{+}23]. hypoxia [BAL^{+}22, LDJL^{+}24].

ichthyoplankton [NYN^{+}22]. iconic [Kit22]. identification [IM22]. identifying [HMS^{+}23]. illucens [LSU^{+}22]. image [TAL^{+}22]. immersion [AKH^{+}23]. immune [DFG22]. Impacts [DFG22, MMG23, ADY^{+}22, MNI^{+}22]. Implications [Pla22, ACA22]. importance [AKK^{+}23]. impoundment [MMM21]. improved [MKPL23, MBT^{+}22, DCM^{+}23]. improvements [KTT22]. improves [LSU^{+}22, OOMM23, ZOGO22]. Improving [CMN^{+}23]. in-river [Los22]. including [AOR^{+}22]. increase [Los22]. Increased [ASS^{+}23]. India [BDM21]. indigenous [OMO^{+}23]. individualised [TAL^{+}22]. Indonesia [BDM21]. induces [UPSP22]. industry [DCM^{+}23]. Influence [AOR^{+}22, KSH23, MKZ^{+}21]. influences [CMR23, FFH^{+}22]. influencing [SEN23]. ingredient [HHR^{+}23]. inhibens [TLS^{+}23]. inland [MMM21]. innovations [MBD^{+}23]. Insights [KNM^{+}23]. Integration [MBT^{+}22, ZOB^{+}23]. intensification [SMD^{+}22]. intensive [ARJ^{+}22, WNS^{+}22]. intensively [FFH^{+}22]. intertidal [KTT22]. intestinal [HMS^{+}23, NHP^{+}22]. intestine [IHA^{+}22]. introgression [PGT^{+}23]. invasion [RCC^{+}23]. invasive [APT^{+}22, BAL^{+}22]. invertebrates [BAL^{+}22]. Ipomoea [RZA^{+}22]. iron [CMW^{+}22]. Islands [TDT23]. isolated [DJF^{+}22].

J [Ano23b]. Japan [Kit22]. Japanese [NIYK22]. japonicus [KTT22, NIYK22]. Jasus [DFG22]. juvenile [AKH^{+}23, IHA^{+}22, Pla22, ZJS^{+}21]. juveniles [GRJM^{+}23, HPPL23].

Kariba [MMG23, MNN^{+}22]. Kenya [MNI^{+}22, OMOO22]. Kenyan [OMO^{+}23]. keratitis [MEHS22]. key [BDGS21]. knowledge [AKK^{+}23, Kit22, MVN^{+}24, OMO^{+}23, PC21].

L [BSH^{+}22]. L-methionine [BSH^{+}22]. L. [AMS^{+}22, LSU^{+}22, SMM^{+}22]. Labeo [ASSR21]. laboratory [NG23]. laboratory-reared [NG23]. lactate [PGT^{+}23]. Lactuca [SMM^{+}22]. lagocephalus [PC21]. Lake [MKPL23, MMG23, OMO^{+}23, BAL^{+}22, FSS^{+}23, NG23, BAL^{+}22, MNN^{+}22, MVN^{+}24, OMOO22]. Lamarck [OYO22]. Lanka [DAJ^{+}23]. large [ASS^{+}23]. largest [MMM21]. larvae [LSU^{+}22, NIYK22]. larval [CMR23, FFH^{+}22]. larviculture [KOK^{+}22]. Lawrence [GFL22]. leader [Los22]. length [Los22]. Lethrinus [TDT23]. lettuce [SMM^{+}22]. lettuce-Nile [SMM^{+}22]. levels [ASS^{+}23]. Life [TDT23, HP22, RSAM23, UOCT23]. life-cycle [HP22]. life-stages [RSAM23]. limited [BAL^{+}22, OOMM23]. Linnaeus [MNI^{+}22, SPZT23, ZJS^{+}21]. lipped [UvdRJ23]. literature [HMS^{+}23]. Litopenaeus [NHP^{+}22, WNS^{+}22]. live [HPPL23, KOK^{+}22, DCM^{+}23]. livelihoods [MNI^{+}22]. liver [IHA^{+}22, RBW^{+}22]. Lobsters [DFG22]. Local [AKK^{+}23, MNI^{+}22]. Locomotory [KSH23]. Long [Kit22]. Long-term [Kit22]. longlines [ZSR23]. Lowlands [RCC^{+}23]. lucioperca [MKZ^{+}21]. lucius [GFL22]. lupus [LDJL^{+}24].

macroalgae [SJ23]. macrobrachion [KNM^{+}23]. Macrobrachium [DAJ^{+}23, DJF^{+}22, IIB^{+}23, KNM^{+}23]. macroscopic [HBP22]. major [NYN^{+}22]. male [AMS^{+}22, MKZ^{+}21]. Man [DAJ^{+}23]. management [CMW^{+}22, MMG^{+}22, OMO^{+}23, TLS^{+}23]. mangrove [OYO22]. mapping [RSAM23]. Mariana [TDT23]. mariculture [LMD^{+}23]. marine [HMK^{+}22]. massoniana [AMS^{+}22]. maturation [MKZ^{+}21]. meal [BSH^{+}22, HHR^{+}23, IM22, LSU^{+}22]. meals [KMMM23]. measure [TAL^{+}22]. measurement [DCM^{+}23]. Mekong [SMD^{+}22]. melanostictus [NIYK22]. Mercenaria [HPPL23]. meristic [MNN^{+}22]. metabarcoding [IM22]. metacercaria [USB^{+}22]. methionine [BSH^{+}22]. method [RPM^{+}23]. methods [MNN^{+}22, OYO22, RSAM23]. microalgae [HPPL23]. microalgal [KOK^{+}22]. microbial [GFL22]. microbiome [HMS^{+}23]. Microbiota [MLHLGV^{+}23, GFL22]. Mid [BWM22]. Mid-Atlantic [BWM22]. mimicry [AOR^{+}22]. minor [LDJL^{+}24, LDJL^{+}24]. mitigating [ADY^{+}22]. mixed [MEHS22]. model [EAAGCMR23, HP22, RPM^{+}23]. Modelling [NF22, RSAM23]. moderate [LDJL^{+}24]. monitoring [LL23]. monodon [ARJ^{+}22]. monosex [AMS^{+}22]. Mori [KTT22]. morpho [SBJJ22]. morpho-physiological [SBJJ22]. morphological [MNN^{+}22]. mortality [BWM22, CMW^{+}22, RPM^{+}23, ZJS^{+}21]. mossambicus [UPSP22]. Moxos [RCC^{+}23]. mud [SAW^{+}22]. Mukosi [MMM21]. mullet [APT^{+}22]. multicriteria [MKPL23]. multifactorial [ADY^{+}22]. multistage [HP22]. mussel [BAD^{+}22, UvdRJ23]. mussels [SJ23]. Mwanza [MKPL23]. mykiss [CMW^{+}22, HMS^{+}23].

N [Ano23b]. National [ZOB^{+}23]. natural [BWM22]. net [RPM^{+}23]. newly [MMG^{+}22]. Nile [AMS^{+}22, KMMM23, LSU^{+}22, MBT^{+}22, MNI^{+}22, SMM^{+}22, SPZT23, ZOGO22]. niloticus [AKJ^{+}23, AMS^{+}22, KMMM23, LSU^{+}22, MNN^{+}22, MBT^{+}22, MNI^{+}22, SMM^{+}22, SPZT23, ZOGO22]. nitrite [DJF^{+}22]. noise [ASS^{+}23]. non [BAL^{+}22, KOK^{+}22]. non-microalgal [KOK^{+}22]. non-target [BAL^{+}22]. north [FSS^{+}23]. north-temperate [FSS^{+}23]. Northern [GFL22, TDT23, USB^{+}22]. note [ZSR23]. Notes [LMD^{+}23]. nursery [GFL22, GRJM^{+}23, NIYK22]. nutrient [MBT^{+}22]. Nutrients [KMMM23].

obscura [AKK^{+}23]. ocean [GRJM^{+}23, UOCT23]. offshore [BAD^{+}22]. oilseed [KMMM23]. olivacea [SAW^{+}22]. Ompok [IHA^{+}22]. On-farm [AKJ^{+}23]. Oncorhynchus [CMW^{+}22, HMS^{+}23, Los22, TAL^{+}22]. open [LMD^{+}23, ZSR23]. optimal [DFG22]. Oreochromis [AKJ^{+}23, AMS^{+}22, KMMM23, LSU^{+}22, MNN^{+}22, MBT^{+}22, MNI^{+}22, SMM^{+}22, SPZT23, UPSP22, ZOGO22]. origin [Los22]. Ostrea [ZJS^{+}21]. Overlap [NIYK22]. oxidative [LDJL^{+}24]. oxidizing [CMW^{+}22]. Oxygymnocypris [DTY^{+}21]. oyster [OYO22, ZJS^{+}21].

Pacific [NHP^{+}22, TDT23, WNS^{+}22]. papain [KMMM23]. Parachanna [AKK^{+}23]. parameters [KNM^{+}23]. parasites [USB^{+}22]. part [MBD^{+}23]. partial [HHR^{+}23, IIB^{+}23]. participation [SEN23]. patchiness [Pla22]. pathogenicity [USB^{+}22]. pattern [KSH23]. pearl [CMN^{+}23]. pekinensis [MBT^{+}22]. pelagic [NYN^{+}22]. Penaeus [ARJ^{+}22]. Perca [FSS^{+}23]. perch [CMN^{+}23, FSS^{+}23, RDH^{+}22]. performance [AKJ^{+}23, ACA22, ARJ^{+}22, IHA^{+}22, IIB^{+}23, KMMM23, LSU^{+}22, NHP^{+}22, OYO22, SMM^{+}22, ZOGO22]. performances [RDH^{+}22]. Perinereis [HMK^{+}22]. perspective [ADY^{+}22]. perspectives [ZOB^{+}23]. Phaeobacter [TLS^{+}23]. phase [MKZ^{+}21]. Philippine [LMD^{+}23]. physiological [LDJL^{+}24, SBJJ22]. Pigmentary [MEHS22]. pike [GFL22]. pikeperch [MKZ^{+}21]. pine [AMS^{+}22]. Pinus [AMS^{+}22]. Plan [ZOB^{+}23]. pollen [AMS^{+}22]. polychaete [HMK^{+}22]. ponds [AKJ^{+}23, SPZT23]. Population [SAW^{+}22, BWM22, Kit22, KTT22, KNM^{+}23, MMG23, TAL^{+}22]. populations [AKK^{+}23]. Porifera [LMD^{+}23]. post [DFG22, MMM21]. post-harvest [DFG22]. potential [BAD^{+}22, GRJM^{+}23, KSH23]. practical [HHR^{+}23]. prawn [DAJ^{+}23, IIB^{+}23, KNM^{+}23]. predators [GRJM^{+}23].

Predatory [APT^{+}22]. predict [RSAM23]. preservation [NYN^{+}22]. probiont [TLS^{+}23]. probiotics [KOK^{+}22, KSH23]. procedure [PGT^{+}23]. producers [ADY^{+}22]. producing [AMS^{+}22]. production [MMG^{+}22, MBT^{+}22, MNI^{+}22, WNS^{+}22]. profile [HPPL23, HMK^{+}22]. Profiling [RZA^{+}22]. properties [IHA^{+}22]. protein [NHP^{+}22, ZOGO22]. proteomic [RBW^{+}22]. protocols [DCM^{+}23]. public [ASS^{+}23]. purebreds [RDH^{+}22]. pyrazosulfuron [UPSP22].

quality [DFG22, IIB^{+}23, LL23]. Quantitative [BDGS21]. Queensland [CMN^{+}23].

rainbow [CMW^{+}22, HMS^{+}23]. rapa [MBT^{+}22]. rapid [PGT^{+}23]. reanalysis [Kit22]. reared [AKJ^{+}23, FFH^{+}22, NG23]. recirculating [LL23]. recreationally [Kit22]. recruitment [SEN23]. recycling [MBT^{+}22]. red [PC21]. red-tailed [PC21]. reduce [Los22]. regarding [Ano23b]. region [OMO^{+}23]. reidi [Pla22]. related [ADY^{+}22]. relaying [ZSR23]. release [RPM^{+}23]. reliable [PGT^{+}23]. remote [RBW^{+}22]. replacement [BSH^{+}22]. replacer [HHR^{+}23]. report [MLHLGV^{+}23]. reproductive [HBP22, MMG23, OOMM23]. Republic [ACA22]. research [DCM^{+}23]. reservoir [ADM23, MMM21, MMG^{+}22]. reservoirs [DAJ^{+}23]. resistance [PFM^{+}22]. resource [BDM21, OMO^{+}23]. response [Ano23b]. restoration [AOR^{+}22, BAD^{+}22]. Restrictions [Los22]. retainment [SEN23]. returns [LSU^{+}22]. review [HMS^{+}23, KOK^{+}22, MBD^{+}23, UOCT23]. Revolution [BC21]. Reynolds [Ano23b]. Ribb [MMG^{+}22]. river [Los22, GFL22]. RNAlater [RBW^{+}22]. Rock [DFG22]. role [KOK^{+}22, OMO^{+}23]. rosenbergii [DAJ^{+}23, IIB^{+}23]. rotifer [KSH23]. rotundiformis [KSH23].

S4 [TLS^{+}23]. saithe [ASS^{+}23]. Salinity [CTCH23]. salmon [HP22, Los22, RPM^{+}23, RBW^{+}22, TAL^{+}22]. salmonid [USB^{+}22]. sampling [RBW^{+}22]. Sander [FFH^{+}22, MKZ^{+}21]. sandfish [GRJM^{+}23]. sardine [NIYK22]. Sardinops [NIYK22]. sativa [SMM^{+}22]. saugeye [FFH^{+}22]. scabra [GRJM^{+}23]. scale [ADY^{+}22, SMM^{+}22, SEN23]. scapulare [CMN^{+}23]. sculpins [Ano23b, TE22]. Scylla [SAW^{+}22]. sea [LMD^{+}23, SBJJ22, UOCT23, DCM^{+}23, ZSR23]. seabed [BAD^{+}22]. seahorse [Pla22]. Seale [TDT23]. Seasonal [FSS^{+}23]. Seasonality [USB^{+}22]. sectional [KG23]. sediment [UvdRJ23]. seed [SLU^{+}23, SJ23]. selected [KOK^{+}22]. selection [MKPL23]. semi [WNS^{+}22]. semi-intensive [WNS^{+}22]. separation [SJ23]. sharptooth [SPZT23]. Shellfish [ZSR23]. shift [APT^{+}22]. shock [ADY^{+}22]. shrimp [ARJ^{+}22, BDM21, NHP^{+}22, STF^{+}22, SMD^{+}22, WNS^{+}22]. Sicyopterus [PC21]. signs [Ano23b, TE22]. Simulations [Pla22]. single [EAAGCMR23]. sites [MKPL23]. six [UOCT23]. Size [ERF^{+}23, NYN^{+}22]. small [ADY^{+}22, HBP22, NYN^{+}22, SMM^{+}22, SEN23]. small-scale [ADY^{+}22, SMM^{+}22, SEN23]. smelt [AKH^{+}23, ERF^{+}23, CTCH23]. snakehead [AKK^{+}23]. snapper [UvdRJ23]. socio [ADY^{+}22, MNI^{+}22]. socio-economic [ADY^{+}22, MNI^{+}22]. soft [UvdRJ23]. soft-sediment [UvdRJ23]. soldier [LSU^{+}22]. South [BDGS21]. southern [BWM22, DFG22]. soybean [BSH^{+}22]. sp [LMD^{+}23]. spatial [RSAM23]. Species [IM22, BDGS21, Kit22, MNN^{+}22, NYN^{+}22, RSAM23, UOCT23]. speed [KSH23]. spleen [MLHLGV^{+}23]. splendens [MEHS22]. sponge [LMD^{+}23]. sport [Los22]. spotted [LDJL^{+}24]. spp [USB^{+}22]. Sri [DAJ^{+}23]. St. [GFL22]. stages [RSAM23]. Stakeholder [ZOB^{+}23]. standardised [DCM^{+}23]. State [MVN^{+}24]. statolith [BWM22]. status [BSH^{+}22, DFG22, MBD^{+}23, TDT23]. steroids [AMS^{+}22]. stewartii [DTY^{+}21]. stinging [BSH^{+}22, HHR^{+}23]. Stocking [ZJS^{+}21, SMM^{+}22]. strains [AKJ^{+}23]. stress [LDJL^{+}24]. Strongylocentrotus [SBJJ22]. structure [CTCH23, DTY^{+}21, Kit22, KNM^{+}23, MMG23, NHP^{+}22, TAL^{+}22]. studies [ADM23]. study [ASS^{+}23, KG23, OMO^{+}23, SPZT23]. sturgeon [CMR23, NG23]. Sub [MMG^{+}22, DFG22]. Sub-basin [MMG^{+}22]. sub-optimal [DFG22]. subsistence [ADY^{+}22]. substrates [KTT22]. succession [SEN23]. suitable [MKPL23, NIYK22]. Sundarbans [SAW^{+}22]. sunflower [HHR^{+}23]. supplemental [BSH^{+}22]. supplementation [PFM^{+}22]. support [PC21]. suppress [BAL^{+}22]. Suppression [KTT22]. surveys [NYN^{+}22]. survival [ASSR21, FFH^{+}22, HPPL23, OYO22, PFM^{+}22]. suspension [OYO22]. sustainable [MBD^{+}23]. swimming [KSH23]. SWOT [BDGS21]. synthetic [AMS^{+}22]. System [MVN^{+}24, APT^{+}22, IIB^{+}23, MBD^{+}23, RPM^{+}23, SMM^{+}22, ZOGO22]. Systematic [HMS^{+}23]. systems [ACA22, ARJ^{+}22, LL23].

tailed [PC21]. Tana [MMG^{+}22]. Tanzania [SLU^{+}23]. target [BAL^{+}22, NYN^{+}22]. Technical [SMD^{+}22]. technologies [MNI^{+}22]. technology [MBD^{+}23]. temperate [FSS^{+}23]. temperature [EAAGCMR23, HMK^{+}22, IHA^{+}22]. term [Kit22]. testing [KG23]. testudineus [RDH^{+}22]. Thailand [BDM21]. their [ACA22, SLU^{+}23]. thermal [MKZ^{+}21]. threatened [HP22, Los22]. threats [AKK^{+}23]. Thymallus [AOR^{+}22]. Tiffan [Ano23b, Ano23b]. tigerfish [MMG23]. Tigriopus [KTT22]. tilapia [AKJ^{+}23, AMS^{+}22, KMMM23, LSU^{+}22, MLHLGV^{+}23, MBT^{+}22, MNI^{+}22, SMM^{+}22, SPZT23, UPSP22, ZOGO22]. time [EAAGCMR23]. timing [CTCH23]. tissues [IHA^{+}22]. tool [PC21, TAL^{+}22]. torus [NF22]. total [BSH^{+}22]. traditional [APT^{+}22]. traits [OOMM23]. transcriptomics [SBJJ22]. translocation [Kit22]. transmontanus [CMR23]. transpacificus [ERF^{+}23]. trap [RPM^{+}23]. trauma [Ano23b, TE22]. trials [LMD^{+}23]. triploid [ASSR21]. trout [BAL^{+}22, CMW^{+}22, HMS^{+}23, PGT^{+}23]. tshawytscha [Los22, TAL^{+}22]. Tugwi [MMM21]. tulipa [OYO22]. two [ADM23]. types [ACA22].

Ume [MMG23]. undatum [BWM22]. unfished [BWM22]. urban [IM22]. urchin [SBJJ22, UOCT23]. urchins [DCM^{+}23]. Use [MKPL23, BDM21, FSS^{+}23, MBT^{+}22]. using [AMS^{+}22, BWM22, HHR^{+}23, KOK^{+}22, KTT22, MNN^{+}22, SJ23]. utilisation [BSH^{+}22]. Utilising [WNS^{+}22]. utilization [ZOGO22].

value [SLU^{+}23]. vannamei [NHP^{+}22, WNS^{+}22]. variation [KNM^{+}23]. vegetables [MBT^{+}22]. veronii [PFM^{+}22]. vibriosis [TLS^{+}23]. Victoria [OMOO22, MKPL23, OMO^{+}23]. Vietnam [BDM21, SMD^{+}22]. Vietnamese [RDH^{+}22]. vision [DCM^{+}23]. Vital [AKH^{+}23]. vitreus [FFH^{+}22]. vittatus [MMG23].

warming [UOCT23]. waste [KOK^{+}22]. waste-generated [KOK^{+}22]. Water [LL23, DFG22, GFL22, IIB^{+}23, MBT^{+}22, OOMM23]. weed [RZA^{+}22]. welfare [Pla22]. West [ADY^{+}22, KNM^{+}23, OYO22]. wetland [GFL22]. white [CMR23, NHP^{+}22, WNS^{+}22]. wild [Los22, NG23, SBJJ22]. within [AKK^{+}23, ZOB^{+}23]. wolffish [LDJL^{+}24]. worldwide [UOCT23].

Xestospongia [LMD^{+}23].

year [GFL22]. yellow [FSS^{+}23]. Yellowstone [BAL^{+}22]. yellowtail [TDT23]. young [GFL22]. young-of-year [GFL22]. Youth [SEN23].

Zambia [MNN^{+}22]. Zimbabwe [MMM21, MMG23].

References

Adegbola:2022:FFS

[ACA22] Ygue Patrice Adegbola, Geraud Fabrice Crinot, and Aminou Arouna. Fish farming systems diversity and implications in the Republic of Benin: Fish farm types and their economic performance. Aquaculture, Fish and Fisheries, 2(6):522–539, December 2022. CODEN ???? ISSN 2693-8847.

Agyekumwaa:2023:BES

[ADM23] Elizabeth Agyekumwaa, Sefah Joseph Debrah, and Kwadwo Kesse Mireku. Bio-ecological studies on two cichlids in the Brimsu reservoir of Ghana. Aquaculture, Fish and Fisheries, 3(6):518– 526, December 2023. CODEN ???? ISSN 2693-8847.

Amoussou:2022:MAR

[ADY^{+}22] Toundji Olivier Amoussou, Comlan Eugène Dessouassi, Dorothé Ngondjeb Yong, Siméon Mahougnon Fagnon, Luc Houngbe, Vinsoun Millogo, Ibrahim Imorou Toko, and Emmanuel A. Frimpong. The multifactorial aquaculture-related COVID-19 shock in Benin, West Africa: a socio-economic perspective of mitigating the disruptive impacts on the small-scale and subsistence producers. Aquaculture, Fish and Fisheries, 2 (6):507–521, December 2022. CODEN ???? ISSN 2693-8847.

Afentoulis:2023:VDI

[AKH^{+}23] Virginia Afentoulis, Andrew Kalmbach, Tien-Chieh Hung, Md. Moshiur Rahman, Luke Ellison, and Javier Miranda. Vital dye immersion evaluations with juvenile delta smelt. Aquaculture, Fish and Fisheries, 3(1):102–111, February 2023. CO-DEN ???? ISSN 2693-8847.

Abwao:2023:FGP

[AKJ^{+}23] Jacob Abwao, Domitila Kyule, Joseph O. Junga, James E. Barasa, and Dorcus A. Sigana. On-farm growth performance of different strains of tilapia, Oreochromis niloticus reared in earthen ponds. Aquaculture, Fish and Fisheries, 3(3):247–255, June 2023. CODEN ???? ISSN 2693-8847.

Amoutchi:2023:LFK

[AKK^{+}23] Amien Isaac Amoutchi, Tanoh Marius Kamelan, Alpha Kargbo, Doh Arioste Delchinor Gneho, Essetchi Paul Kouamelan, and Thomas Mehner. Local fishers’ knowledge on the ecology, economic importance, and threats faced by populations of African snakehead fish, Parachanna obscura, within Côte d’Ivoire freshwater ecosystems. Aquaculture, Fish and Fisheries, 3(3):287–301, June 2023. CODEN ???? ISSN 2693-8847.

Aziz:2022:EUP

[AMS^{+}22] Md. Abdul Aziz, Mst. Momtaz Mostary, Israt Jahan Sume, Md. Helal Uddin, Mohd Golam Quader Khan, Md. Samsul Alam, and M. Sadiqul Islam. The efficacy of using pine (Pinus massoniana) pollen as an alternative to synthetic steroids in producing monosex male Nile tilapia (Oreochromis niloticus, L.). Aquaculture, Fish and Fisheries, 2(5):375–383, October 2022. CODEN ???? ISSN 2693-8847.

Anonymous:2023:C

[Ano23a] Anonymous. Corrigendum. Aquaculture, Fish and Fisheries, 3(2):217, April 2023. CODEN ???? ISSN 2693-8847. See [MMG^{+}22].

Anonymous:2023:TRR

[Ano23b] Anonymous. Tiffan response to Reynolds, J. B. 2022. Correspondence regarding: Tiffan, K. F. and N. J. Eller 2022. Backpack electrofishing does not contribute to external signs of gas bubble trauma in sculpins. Aquaculture, Fish and Fish-

eries, 1–6. https://doi.org/10.1002/aff2.81. Aquaculture,

Fish and Fisheries, 3(3):329, June 2023. CODEN ???? ISSN 2693-8847. See [TE22].

Albertson:2022:IBM

[AOR^{+}22] Lindsey K. Albertson, Valerie Ouellet, J. Holden Reinert, Nathan Korb, and Matthew Jaeger. Influence of beaver mimicry restoration on habitat availability for fishes, including Arctic grayling (Thymallus arcticus). Aquaculture, Fish and Fisheries, 2(2):104–115, April 2022. CODEN ???? ISSN 2693-8847.

Akiona:2022:PFD

[APT^{+}22] Anela K. Akiona, Brian N. Popp, Robert J. Toonen, Margaret C. Siple, Keli’i Kotubetey, Hi’ilei Kawelo, and Erik C. Franklin. Predatory fish diets shift towards an invasive mullet in a traditional Hawaiian aquaculture system. Aquaculture, Fish and Fisheries, 2(5):307–320, October 2022. CODEN ???? ISSN 2693-8847.

Ali:2022:EPC

[ARJ^{+}22] Hazrat Ali, Muhammad Meezanur Rahman, Ahmed Jaman, Siddhwartha Kumar Basak, Mahmoud Eltholth, and Francis Murray. Economic performance characterization of intensive shrimp (Penaeus monodon) farming systems in Bangladesh. Aquaculture, Fish and Fisheries, 2(1):57–70, February 2022. CODEN ???? ISSN 2693-8847.

Andersson:2023:INL

[ASS^{+}23] Marica Andersson, Ola Svensson, Terese Swartz, Jack L. Manera, Michael G. Bertram, and Eva-Lotta Blom. Increased noise levels cause behavioural and distributional changes in Atlantic cod and saithe in a large public aquarium — a case study. Aquaculture, Fish and Fisheries, 3(5):447–458, October 2023. CODEN ???? ISSN 2693-8847.

Afroz:2021:GSD

[ASSR21] Khan Benzir Afroz, Md. Saifuddin Shah, Krishna R. Salin, and Md. Lifat Rahi. Growth and survival of diploid and triploid bata, Labeo bata (Hamilton, 1822). Aquaculture, Fish and Fisheries, 1(1):42–50, December 2021. CODEN ???? ISSN 2693-8847.

Bridger:2022:RPO

[BAD^{+}22] Danielle Bridger, Martin J. Attrill, Bede F. R. Davies, Luke A. Holmes, Amy Cartwright, Siân E. Rees, Llucia Mascorda Cabre, and Emma V. Sheehan. The restoration potential of offshore mussel farming on degraded seabed habitat. Aquaculture, Fish and Fisheries, 2(6):437–449, December 2022. CO-DEN ???? ISSN 2693-8847.

Briggs:2022:FCD

[BAL^{+}22] Michelle A. Briggs, Lindsey K. Albertson, Dominique R. Lujan, Lusha M. Tronstad, Hayley C. Glassic, Christopher S. Guy, and Todd M. Koel. Fish carcass deposition to suppress invasive lake trout through hypoxia causes limited, non-target effects on benthic invertebrates in Yellowstone Lake. Aquaculture, Fish and Fisheries, 2(6):470–483, December 2022. CODEN ???? ISSN 2693-8847.

Becker:2021:AFF

[BC21] Joy Becker and Ricardo Calado. Aquaculture, Fish and Fisheries: a new home for the Blue Revolution. Aquaculture, Fish and Fisheries, 1(1):1–2, December 2021. CODEN ???? ISSN 2693-8847.

Babatunde:2021:QSA

[BDGS21] Adeleke Babatunde, Robertson-Andersson Deborah, Moodley Gan, and Taylor Simon. Quantitative SWOT analysis of key aquaculture species in South Africa. Aquaculture, Fish and Fisheries, 1(1):27–41, December 2021. CODEN ???? ISSN 2693-8847.

Boyd:2021:CRU

[BDM21] Claude E. Boyd, Robert P. Davis, and Aaron A. McNevin. Comparison of resource use for farmed shrimp in Ecuador, India, Indonesia, Thailand, and Vietnam. Aquaculture, Fish and Fisheries, 1(1):3–15, December 2021. CODEN ???? ISSN 2693-8847.

Billah:2022:ESM

[BSH^{+}22] Sheikh Masum Billah, Kanij Rukshana Sumi, Sumon Howlader, Subroto Sarkar, Zannatul Ferdous, S. M. Majharul Islam, and Md Shahjahan. Effects of supplemental L-methionine for total replacement of fish meal by soybean meal on growth, feed utilisation and health status of stinging catfish, Heteropneustes fossilis fry. Aquaculture, Fish and Fisheries, 2(5):355–363, October 2022. CODEN ???? ISSN 2693-8847.

Borsetti:2022:ENM

[BWM22] Sarah Borsetti, John Wiedenmann, and Daphne M. Munroe. Evaluation of natural mortality of an unfished gastropod (Buccinum undatum) population using statolith age-frequency data in the southern Mid-Atlantic Bight. Aquaculture, Fish and Fisheries, 2(3):243–251, June 2022. CODEN ???? ISSN 2693-8847.

Campbell:2023:IEG

[CMN^{+}23] Matthew J. Campbell, Mark F. McLennan, Jamie R. Nicolson, Anna Garland, Robert M. Prosser, and Ricky F. Midgley. Improving estimates of growth for pearl perch (Glaucosoma scapulare) in Queensland, Australia. Aquaculture, Fish and Fisheries, 3(1):71–80, February 2023. CODEN ???? ISSN 2693-8847.

Coulter:2023:BID

[CMR23] Angie Coulter, D. Steven O. McAdam, and John S. Richardson. Biotic influences on drift behaviour of larval white sturgeon (Acipenser transmontanus). Aquaculture, Fish and Fisheries, 3(6):550–554, December 2023. CODEN ???? ISSN 2693-8847.

Clift:2022:BMR

[CMW^{+}22] Annie K. Clift, Ashley M. Malmlov, Colby L. Wells, Pete Cadmus, and Paula A. Schaffer. Branchitis and mortality in rainbow trout Oncorhynchus mykiss exposed to iron oxidizing bacteria: Diagnostics and management in a Colorado hatchery. Aquaculture, Fish and Fisheries, 2(3):202–207, June 2022. CO-DEN ???? ISSN 2693-8847.

Chase:2023:SDT

[CTCH23] Samantha N. Chase, Yi-Jiun Jean Tsai, Martin Croshaw, and Tien-Chieh Hung. Salinity and diel timing affect group structure in Delta Smelt. Aquaculture, Fish and Fisheries, 3(5): 407–414, October 2023. CODEN ???? ISSN 2693-8847.

Digamadulla:2023:ACB

[DAJ^{+}23] Dayananda Senathera Digamadulla, Jayasinghe M. Asoka, Clive Jones, Udith K. Jayasinghe-Mudalige, W. M. H. Kelum Wijenayake, Upali S. Amarasinghe, and M. D. S. T. de Croos. Appraisal of culture-based fisheries of giant freshwater prawn (Macrobrachium rosenbergii, De Man, 1879) in reservoirs of Sri Lanka. Aquaculture, Fish and Fisheries, 3(1):81–95, February 2023. CODEN ???? ISSN 2693-8847.

Vos:2023:CCV

[DCM^{+}23] Bas C. De Vos, Mark D. Cyrus, Brett M. Macey, Theodore Batik, and John J. Bolton. Combining computer vision and standardised protocols for improved measurement of live sea urchins for research and industry. Aquaculture, Fish and Fisheries, 3(6):507–517, December 2023. CODEN ???? ISSN 2693-8847.

Day:2022:ISO

[DFG22] Ryan D. Day, Quinn P. Fitzgibbon, and Caleb Gardner. Impacts of sub-optimal water quality on the haemolymph biochemistry, immune status and condition of post-harvest Southern Rock Lobsters (Jasus edwardsii). Aquaculture, Fish and Fisheries, 2(5):321–333, October 2022. CODEN ???? ISSN 2693-8847.

Dutra:2022:GHA

[DJF^{+}22] Fabrício Martins Dutra, Jonatan Raphael Juvenal, Sandra Carla Forneck, Claudia Caramelo Brazão, and Eduardo Luis Cupertino Ballester. Gill and haematological alterations in Macrobrachium amazonicum adults in acute exposure to ammonia and nitrite isolated and combined. Aquaculture, Fish and Fisheries, 2(5):414–424, October 2022. CODEN ???? ISSN 2693-8847.

Dong:2021:GSA

[DTY^{+}21] Le Dong, Guangxiang Tong, Xiaoxing Yang, Lei Li, Ting Yan, Kai Ma, Jiasheng Yin, Bo Ma, and Youyi Kuang. Genetic structure analysis of the cyprinid Oxygymnocypris stewartii. Aquaculture, Fish and Fisheries, 1(1):66–74, December 2021. CODEN ???? ISSN 2693-8847.

Escamilla-Ake:2023:GMF

[EAAGCMR23] Ángel Escamilla-Aké, Luis Enrique Angeles-Gonzalez, Claudia Caamal-Monsreal, and Carlos Rosas. A general model fitting coleoid cephalopod growth as a function of time and temperature to a single curve. Aquaculture, Fish and Fisheries, 3(6): 539–549, December 2023. CODEN ???? ISSN 2693-8847.

Ellison:2023:SFC

[ERF^{+}23] Luke Ellison, Md Moshiur Rahman, Amanda J. Finger, Marade Sandford, Chih-Hsin Hsueh, Andrew A. Schultz, and Tien-Chieh Hung. Size, fecundity and condition factor changes in endangered delta smelt Hypomesus transpacificus over 10 generations in captivity. Aquaculture, Fish and Fisheries, 3(4): 353–365, August 2023. CODEN ???? ISSN 2693-8847.

Fischer:2022:DIS

[FFH^{+}22] Gregory J. Fischer, Tyler J. Firkus, Kendall Holmes, Patrick C. Blaufuss, Jon J. Amberg, and Christopher F. Hartleb. Diet influences survival and growth of intensively reared larval saugeye (Sander vitreus × Sander canadensis). Aquaculture, Fish and Fisheries, 2(6):450–457, December 2022. CODEN ???? ISSN 2693-8847.

Feucht:2023:SHU

[FSS^{+}23] Levi M. Feucht, Logan W. Sikora, Gabrielle P. Shay, Greg G. Sass, and Joseph T. Mrnak. Seasonal habitat use of yellow perch Perca flavescens in a north-temperate lake. Aquaculture, Fish and Fisheries, 3(4):380–387, August 2023. CODEN ???? ISSN 2693-8847.

Gallo:2022:NPE

[GFL22] Benjamin D. Gallo, John M. Farrell, and Brian F. Leydet. Northern pike (Esox lucius) young-of-year gut microbiota and associations with wetland nursery water microbial communities in the St. Lawrence River. Aquaculture, Fish and Fisheries, 2 (5):384–401, October 2022. CODEN ???? ISSN 2693-8847.

Gorospe:2023:CPP

[GRJM^{+}23] Jay R C. Gorospe, Racelle R. Rescordado, Marie Antonette Juinio-Meñez, Margarita dela Torre-dela Cruz, and Paul C. Southgate. Census of potential predators and competitors of sandfish, Holothuria scabra, juveniles during floating hapa ocean nursery culture. Aquaculture, Fish and Fisheries, 3(3): 316–328, June 2023. CODEN ???? ISSN 2693-8847.

Heins:2022:RBS

[HBP22] David C. Heins and Nancy J. Brown-Peterson. The reproductive biology of small fishes and the clutch concept: Combining macroscopic and histological approaches. Aquaculture, Fish and Fisheries, 2(4):253–264, August 2022. CODEN ???? ISSN 2693-8847.

Hossain:2023:EUS

[HHR^{+}23] Anamika Hossain, Md. Amzad Hossain, Md. Golam Rasul, Taslima Akter, Md. Farid Uz Zaman, and Md. Rabiul Islam. Efficacy of using sunflower meal as an ingredient, and partial fishmeal-replacer, in practical feed formulated for stinging catfish (Heteropneustes fossilis). Aquaculture, Fish and Fisheries, 3(3):237–246, June 2023. CODEN ???? ISSN 2693-8847.

Hoang:2022:ETF

[HMK^{+}22] Tung Hoang, Brian Murphy, Lee Chang Kim, Chris Stratford, and Chris Stevenson. Effects of temperature on farmed marine polychaete Perinereis helleri and its fatty acid profile. Aquaculture, Fish and Fisheries, 2(3):216–223, June 2022. CODEN ???? ISSN 2693-8847.

Hines:2023:SLR

[HMS^{+}23] Ian S. Hines, Maggie A. Marshall, Stephen A. Smith, David D. Kuhn, and Ann M. Stevens. Systematic literature review identifying bacterial constituents in the core intestinal microbiome of rainbow trout (Oncorhynchus mykiss). Aquaculture, Fish and Fisheries, 3(5):393–406, October 2023. CODEN ???? ISSN 2693-8847.

Hinrichsen:2022:MGL

[HP22] Richard A. Hinrichsen and Charles M. Paulsen. A multistage Gompertz life-cycle model applied to threatened Chinook salmon. Aquaculture, Fish and Fisheries, 2(1):44–56, February 2022. CODEN ???? ISSN 2693-8847.

Hassan:2023:GSF

[HPPL23] Md Mahbubul Hassan, Edward Perri, Victoria Parks, and Susan Laramore. Growth, survival and fatty acid profile of hard clam, Mercenaria mercenaria, juveniles fed live microalgae diets. Aquaculture, Fish and Fisheries, 3(1):51–60, February 2023. CODEN ???? ISSN 2693-8847.

Islam:2022:ETG

[IHA^{+}22] Md. Rashedul Islam, Mohammad Amzad Hossain, Farjana Afrose, Nirmal Chandra Roy, and Mohammed Mahbub Iqbal. Effect of temperature on the growth performance, haematological properties and histomorphology of gill, intestine and liver tissues in juvenile butter catfish Ompok bimaculatus. Aquaculture, Fish and Fisheries, 2(4):277–286, August 2022. CODEN ???? ISSN 2693-8847.

Islam:2023:EWQ

[IIB^{+}23] Md. Amirul Islam, Shikder Saiful Islam, Joyanta Bir, Prosenjit Debnath, Md. Rahamat Ullah, and Khandaker Anisul Huq. Effect on water quality, growth performance and economics of giant freshwater prawn, Macrobrachium rosenbergii with partial feed in biofloc system. Aquaculture, Fish and Fisheries, 3 (5):435–446, October 2023. CODEN ???? ISSN 2693-8847.

Ido:2022:SIF

[IM22] Atsushi Ido and Takeshi Miura. Species identification in fish meal from urban fisheries biomass with DNA metabarcoding analysis. Aquaculture, Fish and Fisheries, 2(6):562–571, December 2022. CODEN ???? ISSN 2693-8847.

Kyei-Gyamfi:2023:HTA

[KG23] Sylvester Kyei-Gyamfi. HIV testing attitudes among fishers of Elmina fishing community in Ghana: a cross-sectional study. Aquaculture, Fish and Fisheries, 3(6):527–538, December 2023. CODEN ???? ISSN 2693-8847.

Kitada:2022:LTT

[Kit22] Shuichi Kitada. Long-term translocation explains population genetic structure of a recreationally fished iconic species in Japan: Combining current knowledge with reanalysis. Aquaculture, Fish and Fisheries, 2(2):130–145, April 2022. CODEN ???? ISSN 2693-8847.

Kirimi:2023:NDG

[KMMM23] James G. Kirimi, Levi M. Musalia, Adiel Magana, and Jonathan M. Munguti. Nutrients digestibility and growth performance of Nile tilapia (Oreochromis niloticus) fed on oilseed meals with crude papain enzyme. Aquaculture, Fish and Fisheries, 3(1):23–34, February 2023. CODEN ???? ISSN 2693-8847.

Koussovi:2023:GVS

[KNM^{+}23] Guillaume Koussovi, Farokh Niass, Simon Ahouansou Montcho, Clément Agossou Bonou, and Elie Montchowui. Geographical variation of the structure and population parameters of the freshwater prawn Macrobrachium macrobrachion (Herklots, 1851) in West African coastal hydrosystems: Insights from Benin. Aquaculture, Fish and Fisheries, 3(3):302–315, June 2023. CODEN ???? ISSN 2693-8847.

Kagali:2022:CLF

[KOK^{+}22] Robert Nesta Kagali, Erick Ochieng Ogello, Catherine Wachera Kiama, Hee-Jin Kim, Stenly Wullur, Yoshitaka Sakakura, and Atsushi Hagiwara. Culturing live foods for fish larviculture using non-microalgal diet: the role of waste-generated bacteria and selected commercial probiotics — a review. Aquaculture, Fish and Fisheries, 2(2):71–81, April 2022. CODEN ???? ISSN 2693-8847.

Kagali:2023:LBE

[KSH23] Robert Nesta Kagali, Yoshitaka Sakakura, and Atsushi Hagiwara. Locomotory behaviour of euryhaline rotifer Brachionus rotundiformis: the potential influence of probiotics on swimming pattern and speed. Aquaculture, Fish and Fisheries, 3 (6):497–506, December 2023. CODEN ???? ISSN 2693-8847.

Koga:2022:SCI

[KTT22] Shinichi Koga, Yoshiki Takayama, and Tatsuki Toda. Suppression of cannibalism in the intertidal copepod Tigriopus japonicus (Mori, 1932) and improvements in population density using artificial substrates. Aquaculture, Fish and Fisheries, 2(2): 146–150, April 2022. CODEN ???? ISSN 2693-8847.

Francois:2024:DCE

[LDJL^{+}24] N. R. Le François, C. Drouin-Johnson, F. Larouche, D. Chabot, and P. U. Blier. Does chronic exposure to moderate hypoxia affect growth, physiological and oxidative stress biomarkers in adult spotted wolffish (Anarhichas minor) and the hybrid A. minor × A. lupus? Aquaculture, Fish and Fisheries, 4(1):1–13, February 2024. CODEN ???? ISSN 2693-8847.

Lindholm-Lehto:2023:WQM

[LL23] Petra Lindholm-Lehto. Water quality monitoring in recirculating aquaculture systems. Aquaculture, Fish and Fisheries, 3(2):113–131, April 2023. CODEN ???? ISSN 2693-8847.

Lalas:2023:NDE

[LMD^{+}23] Jue Alef A. Lalas, Geminne G. Manzano, Lee Arraby B. Desabelle, Lilibeth A. Salvador-Reyes, Porfirio M. Aliño, and Maria Vanessa Baria-Rodriguez. Notes on the distribution, ecology and open sea mariculture trials of the Philippine blue sponge, Xestospongia sp. (Porifera). Aquaculture, Fish and Fisheries, 3(2):132–148, April 2023. CODEN ???? ISSN 2693-8847.

Losee:2022:RLL

[Los22] James P. Losee. Restrictions on leader length in an in-river sport fishery reduce catch of threatened wild Chinook salmon Oncorhynchus tshawytscha and increase harvest of hatchery origin Chinook. Aquaculture, Fish and Fisheries, 2(4):296–305, August 2022. CODEN ???? ISSN 2693-8847.

Limbu:2022:BSF

[LSU^{+}22] Samwel Mchele Limbu, Amon Paul Shoko, Eusebia Ernest Ulotu, Siwema Amran Luvanga, Fridah Mukiri Munyi, John Obedy John, and Mary Adhiambo Opiyo. Black soldier fly (Hermetia illucens, L.) larvae meal improves growth performance, feed efficiency and economic returns of Nile tilapia (Oreochromis niloticus, L.) fry. Aquaculture, Fish and Fisheries, 2(3):167–178, June 2022. CODEN ???? ISSN 2693-8847.

McCusker:2023:BTP

[MBD^{+}23] Stephen McCusker, Majbritt Bolton Warberg, Simon J. Davies, Cecilia de Souza Valente, Mark P. Johnson, Ronan Cooney, and Alex H. L. Wan. Biofloc technology as part of a sustainable aquaculture system: a review on the status and innovations for its expansion. Aquaculture, Fish and Fisheries, 3(4):331–352, August 2023. CODEN ???? ISSN 2693-8847.

Mulokozi:2022:INT

[MBT^{+}22] Deogratias Pius Mulokozi, Håkan Berg, Rashid Adam Tamatamah, Torbjörn Lundh, and Paul Ochieng Onyango.

Integration of Nile tilapia (Oreochromis niloticus) and vegetables (Amaranthus hybridus and Brassica rapa pekinensis) for improved production, water use efficiency and nutrient recycling. Aquaculture, Fish and Fisheries, 2(6):493–506, December 2022. CODEN ???? ISSN 2693-8847.

McLaughlin:2022:PKM

[MEHS22] Alicia McLaughlin, Ashley Emanuele, Catherine Hadfield, and Kevin Snekvik. Pigmentary keratitis and mixed chromatophoroma in a betta fish (Betta splendens). Aquaculture, Fish and Fisheries, 2(6):587–592, December 2022. CODEN ???? ISSN 2693-8847.

Mabula:2023:UGB

[MKPL23] Makemie Jumanne Mabula, Danielson Kisanga, Siajali Pamba, and Samwel Mchele Limbu. Use of GIS-based multicriteria evaluation for improved selection of suitable sites for cage fish farming in Mwanza Gulf, Lake Victoria. Aquaculture, Fish and Fisheries, 3(6):472–486, December 2023. CODEN ???? ISSN 2693-8847.

Milla:2021:DCP

[MKZ^{+}21] Sylvain Milla, Amine Khendek, Daniel Zarski, Yannick Ledoré, Imen Ben Ammar, and Pascal Fontaine. Duration of chilling phase, but not thermal condition, influence the gonad maturation of male and female domesticated pikeperch (Sander lucioperca). Aquaculture, Fish and Fisheries, 1(1):51–59, December 2021. CODEN ???? ISSN 2693-8847.

Martinez-Lara:2023:MAE

[MLHLGV^{+}23] Pablo Martínez-Lara, Jorge Hernández-López, Estefanía Garibay-Valdez, Diana Medina-Félix, Marcel Martínez-Porchas, Daniel Coronado-Molina, Romel J. Ortiz-Luna, José H. Puerto, and Martina Hilda Gracia-Valenzuela. Microbiota attached to and encapsulated by granulomas dissected from tilapia spleen: a case report. Aquaculture, Fish and Fisheries, 3(1):96–101, February 2023. CODEN ???? ISSN 2693-8847.

Mequanent:2022:NBR

[MMG^{+}22] Dagnew Mequanent, Minwyelet Mingist, Abebe Getahun, Wassie Anteneh, and Benyam Hailu. The newly built Ribb Reservoir fisheries, Tana Sub-basin, Ethiopia: New fishery establishment, diversity, production, challenges and management. Aquaculture, Fish and Fisheries, 2(3):189–201, June 2022. CODEN ???? ISSN 2693-8847. See correction [Ano23a].

Magqina:2023:IAF

[MMG23] Terence Magqina, Ashley Mudungwe, and David Goza. Impacts of artisanal fishing on the reproductive biology and population structure of tigerfish (Hydrocynus vittatus) in the Ume Basin of Lake Kariba, Zimbabwe. Aquaculture, Fish and Fisheries, 3(5):415–423, October 2023. CODEN ???? ISSN 2693-8847.

Magqina:2021:FDC

[MMM21] Terence Magqina, Chipo Mungenge, and Kuzivakwashe A. Mawoyo. Fish diversity and composition of Tugwi Mukosi Dam, Zimbabwe’s largest inland reservoir post impoundment. Aquaculture, Fish and Fisheries, 1(1):75–83, December 2021. CO-DEN ???? ISSN 2693-8847.

Munguti:2022:NTO

[MNI^{+}22] Jonathan M. Munguti, Rita Nairuti, Jacob O. Iteba, Kevin O. Obiero, Domitila Kyule, Mary A. Opiyo, Jacob Abwao, James G. Kirimi, Nicholas Outa, Mavindu Muthoka, Cecilia M Githukia, and Erick O. Ogello. Nile tilapia (Oreochromis niloticus Linnaeus, 1758) culture in Kenya: Emerging production technologies and socio-economic impacts on local livelihoods. Aquaculture, Fish and Fisheries, 2(4):265–276, August 2022. CODEN ???? ISSN 2693-8847.

Makeche:2022:CON

[MNN^{+}22] Mauris Chinyama Makeche, Tamuka Nhiwatiwa, Joseph Ndebe, Malala Mulavu, Cynthia Sipho Khumalo, Edgar Simulundu, Katendi Changula, Simba Chitanga, Benjamin Mubemba, and Walter Muleya. Characterisation of Oreochromis niloticus fish species of Lake Kariba, Zambia, using morphological, meristic and genetic methods. Aquaculture, Fish and Fisheries, 2(2):116–129, April 2022. CODEN ???? ISSN 2693-8847.

Musinguzi:2024:SKA

[MVN^{+}24] Laban Musinguzi, Nathan Vranken, Vianny Natugonza, William Okello, Maarten van Steenberge, and Jos Snoeks. State of knowledge of aquatic ecosystem and fisheries of the Lake Edward System, East Africa. Aquaculture, Fish and Fisheries, 4(1):1–25, February 2024. CODEN ???? ISSN 2693-8847.

Nishikawa:2022:MTB

[NF22] Haruka Nishikawa and Hisashi Fukushima. Modelling torus behaviour in a fish farming cage. Aquaculture, Fish and Fisheries, 2(3):151–166, June 2022. CODEN ???? ISSN 2693-8847.

Naznin:2023:CAG

[NG23] Afroza Naznin and Janet Genz. Composition analysis of gastrointestinal contents of laboratory-reared and wild lake sturgeon (Acipenser fulvescens). Aquaculture, Fish and Fisheries, 3(2):184–195, April 2023. CODEN ???? ISSN 2693-8847.

Novriadi:2022:EFC

[NHP^{+}22] Romi Novriadi, Vivi Endar Herawati, Slamet Budi Prayitno, Seto Windarto, Keith Mertz, and Hoa Nguyen Duy. Effect of fermented corn protein concentrate on growth performance, haemocyte counts, histological structure of hepatopancreas and intestinal condition of Pacific white shrimp Litopenaeus vannamei. Aquaculture, Fish and Fisheries, 2(2):82–93, April 2022. CODEN ???? ISSN 2693-8847.

Nishikawa:2022:OBS

[NIYK22] Haruka Nishikawa, Sachihiko Itoh, Ichiro Yasuda, and Kosei Komatsu. Overlap between suitable nursery grounds for Japanese anchovy (Engraulis japonicus) and Japanese sardine (Sardinops melanostictus) larvae. Aquaculture, Fish and Fisheries, 2(3):179–188, June 2022. CODEN ???? ISSN 2693-8847.

Nyuji:2022:EFP

[NYN^{+}22] Mitsuo Nyuji, Michio Yoneda, Masahiro Nakamura, Mikio Watai, Tohya Yasuda, and Akinori Takasuka. Effects of formalin preservation on egg size of small pelagic fish as major target species in ichthyoplankton surveys. Aquaculture, Fish and Fisheries, 2(3):208–215, June 2022. CODEN ???? ISSN 2693-8847.

Obiero:2023:RIK

[OMO^{+}23] Kevin Odhiambo Obiero, Jimmy Brian Mboya, Kevin Okoth Ouko, Elijah Migiro Kembenya, Elizabeth Akinyi Nyauchi, Jonathan Mbonge Munguti, Nicholas Otieno Outa, and Cecilia Muthoni Githukia. The role of indigenous knowledge in fisheries resource management for aquaculture development: a case study of the Kenyan Lake Victoria region. Aquaculture, Fish and Fisheries, 3(2):175–183, April 2023. CODEN ???? ISSN 2693-8847.

Obiero:2022:EFF

[OMOO22] Kevin Obiero, Jimmy Brian Mboya, Kevin Okoth Ouko, and Dave Okech. Economic feasibility of fish cage culture in Lake Victoria, Kenya. Aquaculture, Fish and Fisheries, 2(6):484– 492, December 2022. CODEN ???? ISSN 2693-8847.

Ogello:2023:BGW

[OOMM23] Erick O. Ogello, Nicholas O. Outa, Bramwel O. Mukaburu, and Mavindu Muthoka. Biofloc and green water condition improves reproductive traits and fatty acid composition of Artemia franciscana cultured under limited algal conditions. Aquaculture, Fish and Fisheries, 3(1):61–70, February 2023. CODEN ???? ISSN 2693-8847.

Osei:2022:CAG

[OYO22] Isaac Kofi Osei, Kobina Yankson, and Edward Adzesiwor Obodai. Comparative analysis of growth performance and survival of the West African mangrove oyster, Crassostrea tulipa (Lamarck, 1819) cultivated by suspension and bottom culture methods in the Densu Estuary, Ghana. Aquaculture, Fish and Fisheries, 2(3):233–242, June 2022. CODEN ???? ISSN 2693-8847.

Pouil:2021:ATS

[PC21] Simon Pouil and Bérenger Colsoul. Aquaculture as a tool to support goby-fry fishery? Current knowledge on biology and ecology of the red-tailed goby Sicyopterus lagocephalus. Aquaculture, Fish and Fisheries, 1(1):16–26, December 2021. CO-DEN ???? ISSN 2693-8847.

Paul:2022:EDS

[PFM^{+}22] Sagar Paul, Kaniz Farjana, A. G. M. Sofi Uddin Mahamud, Debashis Kumar Mondal, Tamanna Tabassum, Md. Umor Khoiam, Fee Faysal Ahmed, and Tanvir Rahman. Evaluation of the dietary supplementation of autochthonous bacteria on growth, survival and resistance to Aeromonas veronii challenge in Clarias batrachus and Heteropneustes fossilis. Aquaculture, Fish and Fisheries, 2(5):364–374, October 2022. CODEN ???? ISSN 2693-8847.

Padula:2023:RRD

[PGT^{+}23] Anna Padula, Claudia Greco, Lorenzo Talarico, Romolo Caniglia, Caterina Maria Antognazza, Susanna D’Antoni, Massimo Lorenzoni, Isabella Vanetti, Serena Zaccara, and Nadia Mucci. A rapid and reliable detection procedure of Atlantic trout introgression at the diagnostic lactate dehydrogenase chain-1 gene. Aquaculture, Fish and Fisheries, 3(4):388–392, August 2023. CODEN ???? ISSN 2693-8847.

Planas:2022:IHA

[Pla22] Miquel Planas. Implications of holdfast availability and patchiness on juvenile seahorse (Hippocampus reidi) condition and welfare: Simulations in captivity. Aquaculture, Fish and Fisheries, 2(6):552–561, December 2022. CODEN ???? ISSN 2693-8847.

Rusu:2022:RFR

[RBW^{+}22] Anca G. Rusu, James A. Broadbent, Nicholas M. Wade, Cedric J. Simon, Artur N. Rombenso, Simone A. Osborne, and Omar Mendoza-Porras. RNAlater facilitates remote sampling of aquaculture Atlantic salmon liver for proteomic analysis. Aquaculture, Fish and Fisheries, 2(6):578–586, December 2022. CODEN ???? ISSN 2693-8847.

RicoLopez:2023:FBA

[RCC^{+}23] Gabriela Rico Lopez, Claudia Coca Méndez, Joachim Carolsfeld, Oriana Almeida, and Paul Andre Van Damme. Fisheries in a border area of the Moxos Lowlands (Bolivia) after invasion of Arapaima gigas. Aquaculture, Fish and Fisheries, 3(2): 196–210, April 2023. CODEN ???? ISSN 2693-8847.

Rahman:2022:EBG

[RDH^{+}22] Mohammad Lutfar Rahman, Shamima Yeasmin Dipu, Farhana Haque, Sheikh Mohammed Rafiquzzaman, and Mohammad Shafiqul Alam. Evaluation of breeding and growth performances of crossbreds and backcrossbreds with purebreds of Bangladeshi and Vietnamese climbing perch (Anabas testudineus). Aquaculture, Fish and Fisheries, 2(5):343–354, October 2022. CODEN ???? ISSN 2693-8847.

Ruokonen:2023:ETN

[RPM^{+}23] Timo J. Ruokonen, Henni Pulkkinen, Samu Mäntyniemi, Jaakko Erkinaro, and Petri Suuronen. Effect of the trap net emptying method on release mortality of Atlantic salmon estimated by a Bayesian system model. Aquaculture, Fish and Fisheries, 3(4):366–379, August 2023. CODEN ???? ISSN 2693-8847.

Rubec:2023:DMM

[RSAM23] Peter J. Rubec, Christine E. Santi, Jerald S. Ault, and Mark E. Monaco. Development of modelling and mapping methods to predict spatial distributions and abundance of estuarine and coastal fish species life-stages in Florida. Aquaculture, Fish and Fisheries, 3(1):1–22, February 2023. CODEN ???? ISSN 2693-8847.

Roy:2022:PBC

[RZA^{+}22] Sajal Roy, Md. Nazmul Hasan Zilani, Yousof Naser Alrashada, Mohammed Monirul Islam, Fatema Kamrunnaher Akhe, Sk Jamal Uddin, and Md Golam Sarower. Profiling of bioactive compounds and antioxidant activity of aquatic weed Ipomoea aquatica. Aquaculture, Fish and Fisheries, 2(5):425– 435, October 2022. CODEN ???? ISSN 2693-8847.

Sakib:2022:PDM

[SAW^{+}22] Md. Hashmi Sakib, Shawon Ahmmed, Mizanur Rahman Washim, Md. Latiful Islam, and Parvez Chowdhury. Population dynamics of mud crab, Scylla olivacea (Herbst, 1796) from the Sundarbans of Bangladesh. Aquaculture, Fish and Fisheries, 2(3):224–232, June 2022. CODEN ???? ISSN 2693-8847.

Saetra:2022:CMP

[SBJJ22] I. Mathisen Sætra, E. Burgerhout, H. Johnsen, and P. James. Comparative morpho-physiological aspects and transcriptomics of the gonads from wild caught and enhanced green sea urchin (Strongylocentrotus droebachiensis). Aquaculture, Fish and Fisheries, 2(1):28–43, February 2022. CODEN ???? ISSN 2693-8847.

Suh:2023:YRR

[SEN23] Neville N. Suh, Bessy T. Efed, and Richard A. Nyiawung. Youth recruitment and retainment in small-scale fisheries: Factors influencing succession and participation decisions in Cameroon. Aquaculture, Fish and Fisheries, 3(5):424–434, October 2023. CODEN ???? ISSN 2693-8847.

Skelton:2023:ESS

[SJ23] Bradley M. Skelton and Andrew G. Jeffs. Evaluation of separation of seed mussels from macroalgae using chlorination. Aquaculture, Fish and Fisheries, 3(6):487–496, December 2023. CODEN ???? ISSN 2693-8847.

Shoko:2023:FSF

[SLU^{+}23] Amon Paul Shoko, Samwel Mchele Limbu, Eusebia Ernest Ulotu, Salome Daniel Shayo, Mathew Ogalo Silas, Sloans K. Chimatiro, Nazael Amos Madalla, and Rashid Adam Tamatamah. Fish seed and feed value chains analysis and their critical factors for aquaculture development in Tanzania. Aquaculture, Fish and Fisheries, 3(1):35–50, February 2023. CODEN ???? ISSN 2693-8847.

Schuur:2022:TFF

[SMD^{+}22] Anthonie M. Schuur, Aaron A. McNevin, Robert P. Davis, Claude E. Boyd, Susanna Brian, Huynh Quoc Tinh, and Nguyen Phuong Duy. Technical and financial feasibility for intensification of the extensive shrimp farming area in Mekong Delta, Vietnam. Aquaculture, Fish and Fisheries, 2(1):12–27, February 2022. CODEN ???? ISSN 2693-8847.

Sabwa:2022:ESD

[SMM^{+}22] Josiah A. Sabwa, Julius O. Manyala, Frank O. Masese, Kevin Fitzsimmons, Alfred O. Achieng, and Jonathan M. Munguti. Effects of stocking density on the performance of lettuce (Lactuca sativa) in small-scale lettuce-Nile tilapia (Oreochromis niloticus L.) aquaponic system. Aquaculture, Fish and Fisheries, 2(6):458–469, December 2022. CODEN ???? ISSN 2693-8847.

Santi:2023:CSA

[SPZT23] Saïdou Santi, Benjamin Poda, Bilassé Zongo, and Aboubacar Toguyeni. Comparative study of algal community in African sharptooth catfish (Clarias gariepinus, Burchell 1822) and Nile tilapia (Oreochromis niloticus, Linnaeus 1758) farming ponds. Aquaculture, Fish and Fisheries, 3(2):149–164, April 2023. CO-DEN ???? ISSN 2693-8847.

Saborowski:2022:CFA

[STF^{+}22] Reinhard Saborowski, Adrian Tanara, Enno Fricke, Marie Koch, and Wilhelm Hagen. Changes in the fatty acid composition of brown shrimp, Crangon crangon, after boiling. Aquaculture, Fish and Fisheries, 2(5):334–342, October 2022. CODEN ???? ISSN 2693-8847.

Tuckey:2022:AIA

[TAL^{+}22] Nicholas P. L. Tuckey, David T. Ashton, Jiakai Li, Harris T. Lin, Seumas P. Walker, Jane E. Symonds, and Maren Wellenreuther. Automated image analysis as a tool to measure individualised growth and population structure in Chinook salmon (Oncorhynchus tshawytscha). Aquaculture, Fish and Fisheries, 2(5):402–413, October 2022. CODEN ???? ISSN 2693-8847.

Trianni:2023:LHC

[TDT23] Michael S. Trianni, Edward E. DeMartini, and Brett M. Taylor. Life history characteristics and status of the Pacific yellowtail emperor, Lethrinus atkinsoni (Seale 1910), in the Commonwealth of the Northern Mariana Islands. Aquaculture, Fish and Fisheries, 3(2):165–174, April 2023. CODEN ???? ISSN 2693-8847.

Tiffan:2022:BED

[TE22] Kenneth F. Tiffan and Nicole J. Eller. Backpack electrofishing does not contribute to external signs of gas bubble trauma in sculpins. Aquaculture, Fish and Fisheries, 2(6):572–577, December 2022. CODEN ???? ISSN 2693-8847. See response [Ano23b].

Takyi:2023:DEF

[TLS^{+}23] Evelyn Takyi, Jason LaPorte, Saebom Sohn, Rebecca J. Stevick, Erin M. Witkop, Lauren S. Gregg, Amanda Chesler-Poole, Jessica Small, Meredith M. White, Cem Giray, David C. Rowley, David R. Nelson, and Marta Gomez-Chiarri. Development and evaluation of a formulation of probiont Phaeobacter inhibens S4 for the management of vibriosis in bivalve hatcheries. Aquaculture, Fish and Fisheries, 3(3):256–267, June 2023. CODEN ???? ISSN 2693-8847.

Uboldi:2023:HOW

[UOCT23] Thomas Uboldi, Frédéric Olivier, Laurent Chauvaud, and Réjean Tremblay. How ocean warming and acidification affect the life cycle of six worldwide commercialised sea urchin species: a review. Aquaculture, Fish and Fisheries, 3(3):219– 236, June 2023. CODEN ???? ISSN 2693-8847.

Upadhyay:2022:HPE

[UPSP22] Ankur Upadhyay, Parth Pandya, Ankita Salunke, and Pragna Parikh. Herbicide pyrazosulfuron ethyl induces genotoxicity, cell cycle alterations and apoptosis in tilapia (Oreochromis mossambicus). Aquaculture, Fish and Fisheries, 2(4):287–295, August 2022. CODEN ???? ISSN 2693-8847.

Unger:2022:SSP

[USB^{+}22] Patrick Unger, Jaydipbhai Suthar, Falko Baustian, Xaver Neitemeier-Duventester, Sonja Kleinertz, and Harry W. Palm. Seasonality of salmonid parasites from flow-through aquaculture in northern Germany: Emphasis on pathogenicity of Diplostomum spp. metacercaria. Aquaculture, Fish and Fisheries, 2(1):1–11, February 2022. CODEN ???? ISSN 2693-8847.

Underwood:2023:DSC

[UvdRJ23] Lucy H. Underwood, Aimee van der Reis, and Andrew G. Jeffs. Diet of snapper (Chrysophrys auratus) in green-lipped mussel farms and adjacent soft-sediment habitats. Aquaculture, Fish and Fisheries, 3(3):268–286, June 2023. CODEN ???? ISSN 2693-8847.

Walsh:2022:UFE

[WNS^{+}22] Samuel Walsh, Khanh Nguyen, Leila Strebel, Melanie Rhodes, and D. Allen Davis. Utilising feed effectors and automated feeders for semi-intensive Pacific white shrimp (Litopenaeus vannamei) production. Aquaculture, Fish and Fisheries, 2(6): 540–551, December 2022. CODEN ???? ISSN 2693-8847.

Zorita:2021:SDE

[ZJS^{+}21] Izaskun Zorita, Ainhoa Juez, Oihana Solaun, Iñigo Muxika, and José Germán Rodríguez. Stocking density effect on the growth and mortality of juvenile European flat oyster (Ostrea edulis Linnaeus, 1758). Aquaculture, Fish and Fisheries, 1(1): 60–65, December 2021. CODEN ???? ISSN 2693-8847.

Zornu:2023:SPG

[ZOB^{+}23] Jacob Zornu, Matthew Oyih, Martin Binde, Jennifer Viglo, Hayford Agbekpornu, Mary Nkansa, Saraya Tavornpanich, Kari Norheim, Edgar Brun, and Kofitsyo S. Cudjoe. Stakeholder perspectives on the 2023 Ghana National Aquaculture Development Plan: an integration within the ecosystem approach framework. Aquaculture, Fish and Fisheries, 3(6):459– 471, December 2023. CODEN ???? ISSN 2693-8847.

Zablon:2022:BSI

[ZOGO22] Wilfred O. Zablon, Erick O. Ogello, Albert Getabu, and Reuben Omondi. Biofloc system improves protein utilization efficiency and growth performance of Nile tilapia, Oreochromis niloticus fry: Experimental evidence. Aquaculture, Fish and Fisheries, 2(2):94–103, April 2022. CODEN ???? ISSN 2693-8847.

Zorita:2023:SRL

[ZSR23] Izaskun Zorita, Oihana Solaun, and José Germán Rodríguez. Shellfish relaying on longlines in the open sea: a note. Aquaculture, Fish and Fisheries, 3(2):211–216, April 2023. CODEN ???? ISSN 2693-8847.